OpenStudy (anonymous):

x^2 - 40x - 7200 = 0

OpenStudy (anonymous):

@pooja195

OpenStudy (khalilforthewin):

yall got this one i see

OpenStudy (anonymous):

?

OpenStudy (anonymous):

What do you want to solve for @ChiefArnav ?

OpenStudy (anonymous):

x

OpenStudy (anonymous):

Okay, do you know the quadratic formula?

OpenStudy (irishboy123):

no pretty solution: use \(\huge \frac{ -b \pm \sqrt { b^2 - 4 \ a \ c } } { 2a } \)

OpenStudy (anonymous):

Is it 400?

OpenStudy (anonymous):

\[\frac{-(-40)\pm \sqrt{b ^{2}-4(1)(-7200)} }{ 2(1) }\]

OpenStudy (anonymous):

\[\frac{ 40\pm \sqrt{1600 + 28800} }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 40\pm \sqrt{30400} }{ 2 }\]

OpenStudy (anonymous):

\[20\pm\frac{ \sqrt{30400} }{ 2 }\]

OpenStudy (anonymous):

simplify the radical and you'll be done.